OpenStudy (anonymous):

Donna has saved $5,400 for college. She decides to invest some of her money at 3.5% simple interest and the rest at 3% simple interest. In the first year she earned $179 in interest. How many dollars did she earn on the money invested at 3.5% simple interest?

OpenStudy (campbell_st):

well let the 3.5% investment be x the 3% investment is 5400 - x so you know the combined interest is 179 so 179 = 0.03(5400 - x) + 0.035x simplifying and its \[179 = 162 - 0.03x + 0.035x\] which becomes 179 = 162 + 0.005x

OpenStudy (campbell_st):

just solve for x